6/5 = 3m/7 how do I solve for m??
do i cross multiply to get m by itself like this: 6 x 7 = 42 and 3 x 5 =15 so
42 = 15m and 42 divided by 15 is 2.8? I don't think the answer should have a decimal in it.

Answer by Alan3354(69443)

6/5 = 3m/7 how do I solve for m??
-------------
Multiply by 7
42/5 = 3m
Divide by 3
m = 14/5
---------
Or 2.8, same thing.